40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2205     Entry time: Sun Nov 8 22:50:29 2009
Author: Alberto 
Type: Update 
Category: ASC 
Subject: IFO Alignment 

Tonight I aligned the IFO by running the scripts one by one.

SRC was far off and I had to align SRM by hand before the script could work. SPOB is still low when DRM is aligned.

I'm restoring the full IFO now that I'm taking off.

ELOG V3.1.3-